OpenStudy (anonymous):

what is the exact value of radical38 over 2radical6 ?

OpenStudy (anonymous):

over meaning divided by

OpenStudy (anonymous):



OpenStudy (lgbasallote):

\[\large \frac{\sqrt{38}}{2\sqrt 6} \implies \frac{\sqrt{2 \times 19}}{2\sqrt{2 \times 3}} \implies \frac 12 \times \frac{ \sqrt 2}{\sqrt 2} \times \frac{\sqrt{19}}{\sqrt 6} \] do you see it now?

OpenStudy (anonymous):

so would this be the answer?

OpenStudy (anonymous):

@igbasallote

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (lgbasallote):

btw it's LGBASALLOTE